# Transitive Anosov flows on non-compact manifolds

### v1

@@ 3.1. Y-full flows
Fix some small $\varepsilon>0$, let $\eta_{0}$ be an $\varepsilon$-dense periodic orbit of $\varphi$. Then we can write its homology class $[\eta_{0}]=\sum^{n}_{i=1}c_{i}[\alpha_{i}]+[\beta]$, where $[\beta]\in K$ and $c_{i}\in\mathbb{Z}$ for all $i$. If $c_{i}=0$ for all $i$, then $[\eta_{0}]=[\beta]\in K$, so $\eta_{0}$ lifts to an $\varepsilon$-dense union of closed orbits in $Y$, implying that the lifted flow $\varphi_{Y}$ is transitive. So for the rest of this proof, up to renaming, we assume that for some positive integer $k\leq n$, $c_{1},\dots,c_{k}\neq 0$.
Suppose that $c_{1}>0$. By assumption, there exists a periodic orbit that represents the homology class $-[\alpha_{1}]$. By going around that orbit $c_{1}$-times, we obtain a periodic orbit $\gamma_{1}$ that represents the homology class $-c_{1}[\alpha_{1}]$. As in the proof of Theorem B, by applying the Anosov closing lemma and then Fried’s “pair of pants” construction, we find an $\varepsilon$-dense periodic $\eta_{1}$ of $\varphi$ such that $\eta_{1}$ is homologous to $\eta_{0}+\gamma_{1}$. Then
$[\eta_{1}]=[\eta_{0}]+[\gamma_{1}]=\sum^{k}_{i=1}c_{i}[\alpha_{i}]+[\beta]-c_{1}[\alpha_{1}]=\sum^{k}_{i=2}c_{i}[\alpha_{i}]+[\beta].$
If $c_{1}<0$, then by assumption, there exists a periodic orbit that represents the homology class $[\alpha]$, and we obtain $\gamma_{1}$ by going over that orbit $-c_{1}$-times.
@@
$[\eta_{i}]=\sum^{k}_{i_{0}=i+1}c_{i_{0}}[\alpha_{i_{0}}]+[\beta].$
In the end, we obtained an $\varepsilon$-dense periodic orbit $\eta_{k}$ of $\varphi$ such that $[\eta_{k}]=[\beta]\in K$. Then $\eta_{k}$ lifts to an $\varepsilon$-dense union of closed orbits of $\varphi_{Y}$, implying that $\varphi_{Y}$ is transitive, and, by Theorem B, that $\varphi$ is $Y$-full. ∎
@@ 3.2. Manifolds with RFRS fundamental groups and their covers
First, recall that Theorem 1.1 implies that, up to taking a double cover, for any Anosov flow which is not orbit equivalent to a suspension, there exists a pair of distinct periodic orbits $\alpha$, $\alpha^{\prime}$ such that $\alpha$ is freely homotopic to the inverse of $\alpha^{\prime}$. This implies that $[\alpha]$ and $[\alpha^{\prime}]=-[\alpha]$ are both represented by periodic orbits in $H_{1}(M;\mathbb{Z})$. The only issue preventing us from directly applying Theorem 3.2 is that we may in general have $[\alpha]=0\in H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$.
Deep results in $3$-manifolds topology, due to work of Agol and Wise, ensures that for a lot of $3$-manifolds, one can always take finite covers to ensure that $[\alpha]$ is not a torsion element in $H_{1}(M;\mathbb{Z})$. This property, which was introduced by Agol in [Ago08], is that of residually finite rationally solvable (RFRS) groups. We start by recalling its definition:
@@
A group $G$ is RFRS if there is a sequence of subgroups $G=G_{0}>G_{1}>G_{2}>\cdots$ such that $G_{i}$ is a normal subgroup of $G$ for all $i$, $\bigcap_{i}G_{i}=\{1_{G}\}$, $[G\colon G_{i}]<\infty$, and $G_{i+1}\leq(G_{i})^{(1)}_{r}$, where $(G_{i})^{(1)}_{r}=\{x\in G\mid\exists k\neq 0,x^{k}\in[G_{i}\colon G_{i}]\}$.
@@
Let $M$ be an orientable, irreducible $3$-manifold. Suppose that $M$ is not a graph manifold without a non-positively curved metric, then $\pi_{1}(M)$ is virtually special. In particular, $\pi_{1}(M)$ is virtually RFRS.
@@
To prove Theorem A, we will further need the following lemma.
@@
Let $G$ be an RFRS group. Then for any cyclic subgroup $\langle g\rangle$ of $G$, there exists a finite-index subgroup $G^{\prime}$ such that $g\in G^{\prime}$ and such that $g$ represents a nontrivial element $[g]$ in the torsion-free abelian group $H_{1}(G^{\prime};\mathbb{Z})/\mathrm{Torsion}$.
@@
For all $i$, the map $G_{i}\to G_{i}/G_{i+1}$ factors through the map $G_{i}\to H_{1}(G_{i};\mathbb{Z})/\mathrm{Torsion}$. In other words, the following is a commutative diagram
[…]
@@
$H_{1}(G_{i};\mathbb{Z})/\mathrm{Torsion}$ $=\frac{G_{i}/[G_{i},G_{i}]}{\mathrm{Torsion}}$
@@
The last equality follows from the fact that $[G_{i},G_{i}]<(G_{1})^{(1)}_{r}$.
@@
$G_{i}\longrightarrow H_{1}(G_{i};\mathbb{Z})/\mathrm{Torsion}$ $\cong G_{i}/(G_{i})^{(1)}_{r}\longrightarrow\frac{G_{i}/(G_{i})^{(1)}_{r}}{G_{i+1}/(G_{i})^{(1)}_{r}},$
@@
${g}$${g(G_{i})^{(1)}_{r}}$${g(G_{i})^{(1)}_{r}\left(G_{i+1}/(G_{i})^{(1)}_{r}\right).}$
Now, going back to the proof of the lemma, consider the cyclic subgroup $\langle g\rangle<G$. Let $G^{\prime}=G_{i}$, where $G_{i}$ is the subgroup in the normal sequence of subgroups of $G$ such that $g\in G_{i}$ and $g\notin G_{i+1}$. Equivalently, we have that $\langle g\rangle<G_{i}$ and $\langle g\rangle\nless G_{i+1}$. Then the map $G_{i}\to G_{i}/G_{i+1}$ takes $g$ to a nontrivial element in $G_{i}/G_{i+1}$, and it follows from Claim 3.7 that $g$ represents a nontrivial element in $H_{1}(G^{\prime};\mathbb{Z})/\mathrm{Torsion}$. ∎
@@
We will build the desired cover $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ explicitly. We will first find a finite cover $\hat{M}$ of $M$ such that all periodic orbits of $\varphi$ on branching leaves — if there are any — represent nontrivial elements in $H_{1}(\hat{M};\mathbb{Z})/\mathrm{Torsion}$. Then we will construct $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ by applying Theorem 3.2.
@@
Our first step will consist of showing that one can take a regular finite cover $\hat{M}$ of $M$ such that the lifts of (powers of) all the $g_{i}$ represent non-torsion elements in $H_{1}(\hat{M};\mathbb{Z})$.
@@
Consider the subcollection $\{g_{n_{1}},\dots,g_{n_{k}}\}$ of $\{g_{1},\dots,g_{n}\}$ such that $[g_{n_{i}}]=0\in H_{1}(M;Z)/\mathrm{Torsion}$. Up to renaming the $g_{i}$’s, we assume that $g_{1},\dots,g_{k}$ with $k\leq n$ are torsion elements in $H_{1}(M;\mathbb{Z})$. Fix an RFRS sequence for $\pi_{1}(M)$. By Lemma 3.6, we know that for each $g_{i}$, there exists a finite-index normal subgroup $G_{i}$ of $\pi_{1}(M)$ in the fixed RFRS sequence such that $[g_{i}]$ is nontrivial in $H_{1}(G_{i};\mathbb{Z})/\mathrm{Torsion}$. Up to renaming $\{g_{1},g_{2},\dots,g_{k}\}$ again, we assume that $G_{1}$ is the largest subgroup among $\{G_{i}\}_{1\leq i\leq k}$, i.e., $G_{1}$ is the first subgroup appearing in the fixed RFRS sequence of $\pi_{1}(M)$, so that $g_{1},\dots,g_{n}\in G_{1}$. Let $M_{1}$ be the finite cover of $M$ corresponding to $G_{1}$, i.e., $G_{1}=\pi_{1}(M_{1})<\pi_{1}(M)$ (recall that by definition of RFRS, $G_{1}\triangleleft\pi_{1}(M)$ so $M_{1}$ is a regular cover). Then $[g_{1}]$ is nontrivial in $H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$. Note that if $k<n$, i.e., there exists some $g_{i}$ such that $[g_{i}]\neq 0\in H_{1}(M;\mathbb{Z})/\mathrm{Torsion}$, then since $M_{1}$ is a finite cover of $M$, up to replacing $g_{i}$ by some finite power of it, we have $g_{i}\in\pi_{1}(M_{1})$ and $[g_{i}]\neq 0\in H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$.
@@
We also note that it is possible to have other $[g_{i}]$ to be nontrivial in $H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$. This does not pose any issue, as we are able to produce a strictly smaller collection of potentially trivial classes of $H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$ represented by periodic orbits of $\varphi$.
@@
Now, since subgroups of an RFRS group are also RFRS, we can repeat the process for $\pi_{1}(M_{1})$. More precisely, $g_{1},\dots,g_{n}$ are all elements of $\pi_{1}(M_{1})$, and at least $g_{1}$ is such that $[g_{1}]\neq 0\in H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$. Now consider the subcollection $\{g_{n_{1}},\dots,g_{n_{l}}\}$ of $\{g_{2},\dots,g_{n}\}$ such that $[g_{n_{j}}]=0\in H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$. Up to renaming the $g_{i}$, we may assume as above that $g_{2},\dots,g_{l}$, with $l\leq n$ are torsion elements in $H_{1}(M_{1};\mathbb{Z})$. As above, we fix a RFRS sequence for $M_{1}$ to get a subgroup $G_{2}$ (which is finite index in $G_{1}$) such that $g_{2},\dots,g_{l}\in G_{2}$, and, up to further renaming, $[g_{2}]$ is nontrivial in $H_{1}(G_{2};\mathbb{Z})/\mathrm{Torsion}$. Then we consider $M_{2}$ the finite regular cover of $M_{1}$ such that $\pi_{1}(M_{2})=G_{2}$. By construction, in $M_{2}$, $[g_{2}]$ represents a non-torsion element of $H_{1}(M_{2};\mathbb{Z})$. Moreover, since $M_{2}$ is a finite cover of $M_{1}$ and $[g_{1}]\neq 0\in H_{1}(M_{1};\mathbb{Z})/\mathrm{Torsion}$, we deduce that, up to replacing $g_{1}$ by some finite power of it so that $g_{1}\in\pi_{1}(M_{2})$, we have $[g_{1}]\neq 0\in H_{1}(M_{2};\mathbb{Z})/\mathrm{Torsion}$.
This inductive process terminates in at most $n$ steps, and thus we obtained a finite regular cover $\hat{M}$ of $M$, such that appropriate powers of $g_{1},\dots,g_{n}$ all represent non-torsion elements in $H_{1}(\hat{M};\mathbb{Z})$.
@@
Now we have a finite collection of nontrivial elements $[g_{1}],[g_{2}],\dots,[g_{n}]\in H_{1}(\hat{M};\mathbb{Z})/\mathrm{Torsion}$. By construction each $[g_{i}]$ and $[g_{i}^{-1}]=-[g_{i}]$ are represented by periodic orbits of the lifted flow $\hat{\varphi}$ on $\hat{M}$. Moreover, since $\hat{M}$ is a finite cover of $M$, $\hat{\varphi}$ is transitive. Then by applying Theorem 3.2, we obtain an abelian cover $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ of $\hat{M}$ such that the flow $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ obtained by lifting $\hat{\varphi}$ is transitive.
@@
To prove the last statement of the theorem, notice that, by construction, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ does not contain any element conjugated to a power of one of the $g_{i}$. Now, the $g_{i}$ where chosen so that any branching leaf in the orbit space of $\varphi$ (which is the same as the orbit space of $\bar{\varphi}$) was fixed by a conjugate of (any power of) one of the $g_{i}$. Hence, no element in $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ fixes any branching leaf. ∎
@@
It is very likely that for any transitive Anosov flow on a graph-manifold that is not a suspension, one can build a finite cover such that the condition of Theorem 3.2 is satisfied for at least one non-torsion element in $H_{1}(M;\mathbb{Z})$. For instance, if $\phi$ is a transitive Anosov flow on $M$ and there exists an essential torus $T$ in $M$ such that the map $H_{1}(T;\mathbb{Z})\to H_{1}(M;\mathbb{Z})$ is injective, then one can use Theorem 3.2444This is because for any essential torus in a manifold supporting an Anosov flow that is not a suspension, there always exists an element $[\alpha]\in H_{1}(T;\mathbb{Z})$ such that both $[\alpha]$ and $-[\alpha]$ are represented by periodic orbits of the flow. See, e.g., [BF13, Theorem 6.10]. to build an infinite cyclic cover $Y$ on which the lifted flow will be transitive.
@@
A Smale chain is a bi-infinite chain of lozenges $\mathcal{W}=\{L_{i}\}_{i\in\mathbb{Z}}$ that satisfies the following properties:
@@
    each open lozenge is wandering [^5] An open lozenge $L$ is said to be wandering for an Anosov-like action of $G$ if no points of $L$ are fixed by any non-trivial element of $G$..
@@
Theorem 1.19 of [BBM24] gives that if an Anosov-like action is not topologically transitive, then there exists at least one such Smale chain. Equivalently, one can characterize the transitivity of Anosov-like actions as those that do not admit any Smale chain. Note that, while condition (i) in Definition 4.2 is dynamical in nature, conditions (ii) and (iii) only involve the configuration of foliations in the plane, and not the group action. Thus, we can deduce the following sufficient condition for all Anosov-like actions to be transitive on certain bifoliated planes:
@@
Let $\mathcal{P}$ be a bifoliated plane. Assume that $\mathcal{P}$ does not contain any chains of lozenges $\{L_{i}\}_{i\in\mathbb{Z}}$ such that
@@
Let $\varphi\colon M\to M$ be a transitive Anosov flow on a compact 3-manifold $M$. Let $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ be a regular cover of $M$ and $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ be the lift of $\varphi$ to $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$. If $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ has a periodic orbit, then the action $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)\curvearrowright P_{\varphi}$ satisfies Axioms (A1), (A2), and (A3) of an Anosov-like action.

### v2

@@ 3.1. Y-full flows
Fix some small $\varepsilon>0$, let $\eta_{0}$ be an $\varepsilon$-dense periodic orbit of $\varphi$. Then we can write its homology class $\eta_{0}]=\sum^{n}_{i=1}c_{i}[\alpha_{i}]+[\beta]$, where $\beta]\in K$ and $c_{i}\in\mathbb{Z}$ for all $i$. If $c_{i}=0$ for all $i$, then $\eta_{0}]=[\beta]\in K$, so $\eta_{0}$ lifts to an $\varepsilon$-dense union of closed orbits in $Y$, implying that the lifted flow $\varphi_{Y}$ is transitive. So for the rest of this proof, up to renaming, we assume that for some positive integer $k\leq n$, $c_{1},\dots,c_{k}\neq 0$.
Suppose that $c_{1}>0$. By assumption, there exists a periodic orbit that represents the homology class $\alpha_{1}]$. By going around that orbit $c_{1}$-times, we obtain a periodic orbit $\gamma_{1}$ that represents the homology class $c_{1}[\alpha_{1}]$. As in the proof of Theorem B, by applying the Anosov closing lemma and then Fried’s “pair of pants” construction, we find an $\varepsilon$-dense periodic $\eta_{1}$ of $\varphi$ such that $\eta_{1}$ is homologous to $\eta_{0}+\gamma_{1}$. Then
$\eta_{1}]=[\eta_{0}]+[\gamma_{1}]=\sum^{k}_{i=1}c_{i}[\alpha_{i}]+[\beta]-c_{1}[\alpha_{1}]=\sum^{k}_{i=2}c_{i}[\alpha_{i}]+[\beta].$
If $c_{1}<0$, then by assumption, there exists a periodic orbit that represents the homology class $\alpha]$, and we obtain $\gamma_{1}$ by going over that orbit $c_{1}-$times.
@@
$\eta_{i}]=\sum^{k}_{i_{0}=i+1}c_{i_{0}}[\alpha_{i_{0}}]+[\beta].$
In the end, we obtained an $\varepsilon$-dense periodic orbit $\eta_{k}$ of $\varphi$ such that $\eta_{k}]=[\beta]\in K$. Then $\eta_{k}$ lifts to an $\varepsilon$-dense union of closed orbits of $\varphi_{Y}$, implying that $\varphi_{Y}$ is transitive, and, by Theorem B, that $\varphi$ is $Y$-full. ∎
@@ 3.2. Manifolds with RFRS fundamental groups and their covers
First, recall that Theorem 1.1 implies that, up to taking a double cover, for any Anosov flow which is not orbit equivalent to a suspension, there exists a pair of distinct periodic orbits $\alpha$, $\alpha^{\prime}$ such that $\alpha$ is freely homotopic to the inverse of $\alpha^{\prime}$. This implies that $\alpha]$ and $\alpha^{\prime}]=-[\alpha]$ are both represented by periodic orbits. The only issue preventing us from directly applying Theorem 3.3 is that we may in general have $\alpha]=0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$.
Deep results in $3$-manifolds topology, due to work of Agol and Wise, ensures that for a lot of $3$-manifolds, one can always take finite covers to ensure that $\alpha]$ is not a torsion element in $H_{1}(M)$. This property, which was introduced by Agol in [Ago08], is that of residually finite rationally solvable (RFRS) groups. We start by recalling its definition:
@@
A group $G$ is RFRS if there is a sequence of subgroups $G=G_{0}>G_{1}>G_{2}>\cdots$ such that $G_{i}$ is a normal subgroup of $G$ for all $i$, $\bigcap_{i}G_{i}=\{1_{G}\}$, $G\colon G_{i}]<\infty$, and $G_{i+1}\leq(G_{i})^{(1)}_{r}$, where $(G_{i})^{(1)}_{r}=\{x\in G\mid\exists k\neq 0,x^{k}\in[G_{i},G_{i}]\}$.
@@
Let $M$ be an orientable, irreducible $3$-manifold. Suppose that $M$ is not a graph manifold without a non-positively curved metric, then $\pi_{1}(M)$ is virtually special. [^4] In our context, one can define a group to be special if it is a subgroup of a right-angled Artin group, see e.g., [PW18] or [AFW15, Corollary 5.8] for more precisions. In particular, $\pi_{1}(M)$ is virtually RFRS.
@@
To prove Theorem A, we will need the following technical lemma. For a group $G$, we denote by $[G,G]$ its commutator subgroup. and we denote by $H_{1}(G)=H_{1}(G;\mathbb{Z})$ the quotient group $G/[G,G]$.
@@
Let $G$ be an RFRS group. Then for any cyclic subgroup $\langle g\rangle$ (finite or infinite) of $G$, there exists a finite-index subgroup $G^{\prime}$ such that $g\in G^{\prime}$ and such that $g$ represents a nontrivial element $[g]$ in the torsion-free abelian group $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G^{\prime})\mkern-1.5mu}\mkern 1.5mu_{T}$.
@@
For all $i$, the map $G_{i}\to G_{i}/G_{i+1}$ factors through the map $G_{i}\to\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G_{i})\mkern-1.5mu}\mkern 1.5mu_{T}$. In other words, the following is a commutative diagram
$G_{i}$$G_{i}/G_{i+1}$$\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G_{i})\mkern-1.5mu}\mkern 1.5mu_{T}$}
@@
$\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G_{i})\mkern-1.5mu}\mkern 1.5mu_{T}$ $=\frac{G_{i}/[G_{i},G_{i}]}{\mathrm{Tor}(G_{i}/[G_{i},G_{i}])}$
@@
The last equality follows from the fact that $G_{i},G_{i}]<(G_{i})^{(1)}_{r}$.
@@
$G_{i}\longrightarrow\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G_{i})\mkern-1.5mu}\mkern 1.5mu_{T}$ $\cong G_{i}/(G_{i})^{(1)}_{r}\longrightarrow\frac{G_{i}/(G_{i})^{(1)}_{r}}{G_{i+1}/(G_{i})^{(1)}_{r}}$,
@@
[…]
Now, going back to the proof of the lemma, consider the cyclic subgroup $\langle g\rangle<G$. Let $G^{\prime}=G_{i}$, where $G_{i}$ is the subgroup in the RFRS sequence of $G$ such that $g\in G_{i}$ and $g\notin G_{i+1}$. Equivalently, we have that $\langle g\rangle<G_{i}$ and $\langle g\rangle\nless G_{i+1}$. Then the map $G_{i}\to G_{i}/G_{i+1}$ takes $g$ to a nontrivial element in $G_{i}/G_{i+1}$, and it follows from Claim 3.8 that $g$ represents a nontrivial element in $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G^{\prime})\mkern-1.5mu}\mkern 1.5mu_{T}$. ∎
@@
[…]
@@
[…]
@@
Consider the subcollection $\{g_{n_{1}},\dots,g_{n_{k}}\}$ of $\{g_{1},\dots,g_{n}\}$ such that $g_{n_{i}}]=0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$. Up to renaming the $g_{i}$’s, we assume that $g_{1},\dots,g_{k}$ with $k\leq n$ are torsion elements in $H_{1}(M)$. Fix an RFRS sequence for $\pi_{1}(M)$. By Lemma 3.7, we know that for each $g_{i}$, there exists a finite-index normal subgroup $G_{i}$ of $\pi_{1}(M)$ in the fixed RFRS sequence such that $g_{i}]$ is nontrivial in $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G_{i})\mkern-1.5mu}\mkern 1.5mu_{T}$. Up to renaming $\{g_{1},g_{2},\dots,g_{k}\}$ again, we assume that $G_{1}$ is the largest subgroup among $\{G_{i}\}_{1\leq i\leq k}$, i.e., $G_{1}$ is the first subgroup appearing in the fixed RFRS sequence of $\pi_{1}(M)$, so that $g_{1},\dots,g_{n}\in G_{1}$. Let $M_{1}$ be the finite cover of $M$ corresponding to $G_{1}$, i.e., $G_{1}=\pi_{1}(M_{1})<\pi_{1}(M)$ (recall that by definition of RFRS, $G_{1}\triangleleft\pi_{1}(M)$ so $M_{1}$ is a regular cover). Then $g_{1}]$ is nontrivial in $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M_{1})\mkern-1.5mu}\mkern 1.5mu_{T}$. Note that if $k<n$, i.e., there exists some $g_{i}$ such that $g_{i}]\neq 0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$, then since $M_{1}$ is a finite cover of $M$, up to replacing $g_{i}$ by some finite power of it, we have $g_{i}\in\pi_{1}(M_{1})$ and $g_{i}]\neq 0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$.
@@
We also note that it is possible to have other $g_{i}]$ to be nontrivial in $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$. This does not pose any issue, as we are able to produce a strictly smaller collection of potentially trivial classes of $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$ represented by periodic orbits of $\varphi$.
@@
Now, since subgroups of an RFRS group are also RFRS, we can repeat the process for $\pi_{1}(M_{1})$. More precisely, $g_{1},\dots,g_{n}$ are all elements of $\pi_{1}(M_{1})$, and at least $g_{1}$ is such that $g_{1}]\neq 0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$. Now consider the subcollection $\{g_{n_{1}},\dots,g_{n_{l}}\}$ of $\{g_{2},\dots,g_{n}\}$ such that $g_{n_{j}}]=0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M)\mkern-1.5mu}\mkern 1.5mu_{T}$. Up to renaming the $g_{i}$, we may assume as above that $g_{2},\dots,g_{l}$, with $l\leq n$ are torsion elements in $H_{1}(M_{1})$. As above, we fix an RFRS sequence for $M_{1}$ to get a subgroup $G_{2}$ (which is finite index in $G_{1}$) such that $g_{2},\dots,g_{l}\in G_{2}$, and, up to further renaming, $g_{2}]$ is nontrivial in $\mkern 1.5mu\overline{\mkern-1.5muH_{1}(G_{2})\mkern-1.5mu}\mkern 1.5mu_{T}$. Then we consider $M_{2}$ the finite regular cover of $M_{1}$ such that $\pi_{1}(M_{2})=G_{2}$. By construction, in $M_{2}$, $g_{2}]$ represents a non-torsion element of $H_{1}(M_{2})$. Moreover, since $M_{2}$ is a finite cover of $M_{1}$ and $g_{1}]\neq 0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M_{1})\mkern-1.5mu}\mkern 1.5mu_{T}$, we deduce that, up to replacing $g_{1}$ by some finite power of it so that $g_{1}\in\pi_{1}(M_{2})$, we have $g_{1}]\neq 0\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(M_{2})\mkern-1.5mu}\mkern 1.5mu_{T}$.
[…]
@@
Now we have a finite collection of nontrivial elements $g_{1}],[g_{2}],\dots,[g_{n}]\in\mkern 1.5mu\overline{\mkern-1.5muH_{1}(\hat{M})\mkern-1.5mu}\mkern 1.5mu_{T}$. By construction each $g_{i}]$ and $g_{i}^{-1}]=-[g_{i}]$ are represented by periodic orbits of the lifted flow $\hat{\varphi}$ on $\hat{M}$. Moreover, since $\hat{M}$ is a finite cover of $M$, $\hat{\varphi}$ is transitive. Then by applying Theorem 3.3, we obtain an abelian cover $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ of $\hat{M}$ such that the flow $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ obtained by lifting $\hat{\varphi}$ is transitive.
@@
To prove the last statement of the theorem, notice that, by construction, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ does not contain any element conjugated to a power of one of the elements $g_{i}$. Now, the elements $g_{i}$ were chosen so that any branching leaf in the orbit space of $\varphi$ (which is the same as the orbit space of $\bar{\varphi}$) was fixed by a conjugate of (any power of) one of the $g_{i}$. Hence, no element in $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ fixes any branching leaf. ∎
@@
It is very likely that for any transitive Anosov flow on a graph manifold $M$ that is not a suspension, one can build a finite cover such that the condition of Theorem 3.3 is satisfied for at least one non-torsion element in $H_{1}(M)$. For instance, if $M$ supports a transitive Anosov flow and there exists an essential torus $\mathbb{T}$ in $M$ such that the map $H_{1}(\mathbb{T})\to H_{1}(M)$ is injective, then one can use Theorem 3.3 to build an infinite cyclic cover $Y$ on which the lifted flow will be transitive. Indeed, for any essential torus $\mathbb{T}$ in a manifold supporting an Anosov flow that is not a suspension, there always exists an element $\alpha]\in H_{1}(\mathbb{T})$ such that both $\alpha]$ and $\alpha]$ are represented by periodic orbits of the flow; see, e.g., [BF13, Theorem 6.10].
@@
Note that we are now considering all regular covers and not just abelian ones.
@@
Let $G\curvearrowright\mathcal{P}$ be an Anosov-like action. A Smale chain for $G\curvearrowright\mathcal{P}$ is a bi-infinite chain of lozenges $\mathcal{W}=\{L_{i}\}_{i\in\mathbb{Z}}$ such that for all $i$,
@@
    $L_{i}$ and $L_{i+1}$ share a side.
@@
    each lozenge $L_{i}$ is wandering, that is, no points of the interior of $L_{i}$ are fixed by any nontrivial element of $G$.
@@
Theorem 1.19 of [BBM24] gives that if an Anosov-like action is not topologically transitive, then there exists at least one such Smale chain. Equivalently, one can characterize the transitivity of Anosov-like actions as those that do not admit any Smale chain. Note that, while condition (iii) in Definition 4.2 is dynamical in nature, conditions (i) and (ii) only involve the configuration of foliations in the plane, and not the group action. Thus, we can deduce the following sufficient condition for all Anosov-like actions to be transitive on certain bifoliated planes:
@@
Let $\mathcal{P}$ be a bifoliated plane. Assume that $\mathcal{P}$ does not contain any chains of lozenges $\{L_{i}\}_{i\in\mathbb{Z}}$ such that for all $i$,
@@
Let $\varphi\colon M\to M$ be a transitive Anosov flow on a compact 3-manifold $M$. Let $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ be a regular cover of $M$ and $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ be the lift of $\varphi$ to $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$. If $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ has a periodic orbit, then the action $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)\curvearrowright\mathcal{P}_{\varphi}$ satisfies Axioms (A1), (A2), and (A3) of an Anosov-like action.